# TimeSieve: Extracting Temporal Dynamics through Information Bottleneck

### v1

@@ 5. Conclusions
In this paper, we presented the TimeSieve framework for time series forecasting, emphasizing the decomposition and reconstruction of time series data through wavelet and information bottleneck techniques. Our approach integrates Wavelet Decomposition Block (WDB), Wavelet Reconstruction Block (WRB), and Information Fusion and Compression Block (IFCB) to enhance prediction accuracy.
@@
The proposed TimeSieve framework paves the way for future research in time series forecasting, offering a robust and scalable solution that can be applied to diverse domains such as finance, healthcare, and climate science. Future work will focus on further optimizing the model, exploring adaptive wavelet transforms, and extending the framework to multi-variate and multi-modal time series data.
Appendix A Markov Chain Theory
A.1. Basic Concepts
Markov chains are a powerful tool for describing the transitions between states in a system, widely used in various fields such as physics, biology, economics, and computer science. In deep learning and information theory, Markov chains can model the transmission of information and state dependencies. This document will introduce the basic concepts, properties, and applications of Markov chains in information theory.
A Markov chain is a sequence of random variables $\{X_{t}\}_{t\in T}$, where each random variable $X_{t}$ represents the state of the system at time $t$. The core property of a Markov chain is the Markov property, which states that the current state $X_{t}$ depends only on the previous state $X_{t-1}$ and not on any earlier states. Formally, this is expressed as:
$P(X_{t+1}=x_{t+1}\mid X_{t}=x_{t},X_{t-1}=x_{t-1},\ldots,X_{0}=x_{0})$
$\quad=P(X_{t+1}=x_{t+1}\mid X_{t}=x_{t})$
The state space $S$ is the set of all possible states of the Markov chain. For a finite state space, a Markov chain can be represented by a transition matrix $P$, where $P_{ij}$ denotes the probability of transitioning from state $i$ to state $j$:
$P=\begin{bmatrix}P_{11}&P_{12}&\cdots&P_{1n}\\ P_{21}&P_{22}&\cdots&P_{2n}\\ \vdots&\vdots&\ddots&\vdots\\ P_{n1}&P_{n2}&\cdots&P_{nn}\end{bmatrix}$
The transition probabilities must satisfy the following conditions:
1. Non-negativity: $P_{ij}\geq 0$ 2. Normalization: $\sum_{j\in S}P_{ij}=1$
A probability vector $\pi$ is a stationary distribution of a Markov chain if it satisfies $\pi P=\pi$. For an irreducible and aperiodic Markov chain, the stationary distribution exists uniquely, and the state distribution converges to the stationary distribution regardless of the initial distribution:
$\lim_{t\to\infty}P(X_{t}=j)=\pi_{j}$
A.2. Applications in Information Theory
In information theory, Markov chains are commonly used to model the process of information transmission. In particular, in deep learning, hidden layers can be viewed as intermediate states in the information transmission process, where the relationship between input, hidden layer, and output can be modeled as a Markov chain. Assuming the input, hidden layer, and output are denoted as $X$, $Z$, and $Y$ respectively, we have the Markov property $X\rightarrow Z\rightarrow Y$, meaning:
$P(Y\mid Z,X)=P(Y\mid Z)$
Mutual information measures the mutual dependence between two random variables . For the above relationship, mutual information satisfies the following inequality:
$I(X;Z)\geq I(X;Y)$
This indicates that in the process of information transmission, the intermediate layer $Z$ retains more information about the input $X$ than the output $Y$, thereby better predicting the output $Y$.
Appendix B Mutual Information
B.1. Basic Concepts
Mutual Information (MI) is a fundamental concept in information theory that measures the amount of information obtained about one random variable through another random variable. It quantifies the reduction in uncertainty of one variable given knowledge of the other. In the context of deep learning, mutual information can help analyze the dependency and information flow between different layers of a neural network. This section will introduce the basic concepts, properties, and applications of mutual information.
Mutual information between two random variables $X$ and $Y$ is defined as the reduction in the uncertainty of $X$ due to the knowledge of $Y$. Formally, it is given by:
$I(X;Y)=\sum_{x\in\mathcal{X}}\sum_{y\in\mathcal{Y}}p(x,y)\log\left(\frac{p(x,y)}{p(x)p(y)}\right)$
where $p(x,y)$ is the joint probability distribution of $X$ and $Y$, and $p(x)$ and $p(y)$ are the marginal probability distributions of $X$ and $Y$, respectively.
Alternatively, mutual information can be expressed in terms of entropy:
$I(X;Y)=H(X)-H(X\mid Y)=H(Y)-H(Y\mid X)=H(X)+H(Y)-H(X,Y)$
where $H(X)$ and $H(Y)$ are the entropies of $X$ and $Y$, $H(X\mid Y)$ and $H(Y\mid X)$ are the conditional entropies, and $H(X,Y)$ is the joint entropy.
B.2. Properties of Mutual Information
Mutual information has several important properties that make it a useful measure in information theory:
  1. (1)
    Non-negativity: $I(X;Y)\geq 0$. Mutual information is always non-negative.
  2. (2)
    Symmetry: $I(X;Y)=I(Y;X)$. Mutual information is symmetric with respect to $X$ and $Y$.
  3. (3)
    Chain Rule: Mutual information satisfies a chain rule similar to entropy. For three random variables $X$, $Y$, and $Z$:
    $I(X;Y,Z)=I(X;Z)+I(X;Y\mid Z)$
  4. (4)
    Data Processing Inequality: If $X\rightarrow Y\rightarrow Z$ forms a Markov chain, then mutual information satisfies the inequality:
    $I(X;Z)\leq I(X;Y)$
B.3. Applications in Deep Learning
In deep learning, mutual information is used to understand and optimize the information flow through the layers of a neural network. Specifically, the Information Bottleneck (IB) method leverages mutual information to balance compression and prediction. The IB method aims to find a representation $Z$ that retains the most relevant information about the output $Y$ while minimizing the information about the input $X$. Formally, this is expressed as:
$\min\left[I(X;Z)-\beta I(Z;Y)\right]$
where $\beta$ is a trade-off parameter. This approach helps in understanding the trade-offs between compressing the input information and retaining predictive power.
B.4. Mutual Information Estimation
Estimating mutual information from data can be challenging due to the need to estimate the joint and marginal distributions. Various methods have been proposed for mutual information estimation, including:
  1. (1)
    Histogram-based methods: These methods discretize the data into bins and estimate probabilities from the frequency of occurrences.
  2. (2)
    Kernel density estimation: This non-parametric approach uses kernels to estimate the probability density functions.
  3. (3)
    Neural estimation methods: Recent advances use neural networks to approximate mutual information, such as MINE (Mutual Information Neural Estimation).
Appendix C Implementation Details
C.1. Baselines
C.1.1. Koopa (Liu et al., 2023a)
@@ C.1. Baselines
C.1.2. NSTformer (Zeng et al., 2022)
@@ C.2. Datasets
National Illness Dataset: The National Illness dataset, also known as the Influenza-like Illness (ILI) dataset, comprises weekly recorded data from the Centers for Disease Control and Prevention (CDC) of the United States. It covers the period from 2002 to 2021 and includes the ratio of patients exhibiting influenza-like symptoms to the total number of patients seen. This dataset is essential for modeling and forecasting disease outbreaks, aiding in public health planning and response strategies.
Table 4. Summary of Datasets Used in Time Series Forecasting
@@
National Illness 966 Week 7
C.3. Definition of Metric
@@
(18) $\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}\left|y_{i}-\hat{y}_{i}\right|$
@@
(19) $\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-\hat{y}_{i}\right)^{2}$
@@ C.4. Experimental Setup
Table 5. Experimental Setup: Summary of Hardware and Software Configurations
Table 6. Comparison of different $\mathcal{L}_{IB}$ loss weights on Exchange rate dataset. The lookback length is set to $T=2H$.
@@
48 0.139374 0.050472 0.139033 0.045016 0.139010 0.044790 0.138870 0.044050 0.139096 0.043964
96 0.194557 0.110181 0.196048 0.083312 0.195420 0.084286 0.195818 0.085749 0.194953 0.085362
144 0.249113 0.139211 0.243771 0.133058 0.241323 0.116296 0.241018 0.153478 0.240288 0.152741
192 0.300818 0.1886299 0.287500 0.1804375 0.283648 0.1815776 0.283406 0.1777823 0.289262 0.1804006
Table 7. Comparison of different wavelet basis functions on Exchange rate dataset.The lookback length is set to $T=2H$.
@@
48 0.1392049 0.0477886 0.1419515 0.0450065 0.1390332 0.0450161
96 0.1944531 0.1024204 0.1944787 0.0895267 0.1960484 0.0833123
144 0.2446094 0.1230493 0.2429469 0.1325571 0.2437714 0.1330583
192 0.3007113 0.2840106 0.2913505 0.1804375 0.2875000 0.1875995
Figure 3. Comparison of model residuals for TimeSieve, Koopa, LightTS, NSTformer, and Autoformer on the ETTh2 dataset with $H=48$ and $T=96$. The vertical axis represents the absolute difference between predicted and actual values.

### v2

@@ 5 Conclusions
In this paper, we presented the TimeSieve framework for time series forecasting, emphasizing the decomposition and reconstruction of time series data through wavelet and information bottleneck techniques. Our approach integrates Wavelet Decomposition Block, Wavelet Reconstruction Block, and Information Fusion and Compression Block to enhance prediction accuracy.
@@
The proposed TimeSieve framework establishes a foundational approach for time series forecasting across various domains, including finance and climate science. Future work will focus on enhancing the model’s robustness against datasets, alongside exploring extending the model to handle multi-modal time series data.
Appendix A Implementation Details
A.1 Baselines
A.1.1 Koopa [28]
@@
A.1.2 PatchTST [15]
PatchTST is a Transformer-based model designed for efficient and accurate long-term time series forecasting. Developed by researchers at Princeton University and IBM Research, PatchTST introduces two key innovations: the segmentation of time series into subseries-level patches and a channel-independent approach where each channel corresponds to a single univariate time series. These patches serve as input tokens to the Transformer, enabling the model to retain local semantic information, reduce the computational complexity of attention mechanisms, and effectively utilize longer historical sequences.
The model’s architecture leverages the benefits of patching to enhance locality in the embeddings and minimize the quadratic increase in computation and memory usage. The channel-independent design allows PatchTST to achieve significant improvements in forecasting accuracy by sharing the same embedding and Transformer weights across all channels while processing them independently. PatchTST has demonstrated superior performance in both supervised and self-supervised learning tasks, outperforming state-of-the-art Transformer-based models in various benchmark datasets.
A.1.3 DLinear [14]
DLinear is a simple yet effective model designed to challenge the dominance of Transformer-based models in long-term time series forecasting (LTSF). Developed by researchers at the Chinese University of Hong Kong and the International Digital Economy Academy, DLinear is built on the premise that complex architectures like Transformers may not be necessary for capturing temporal dependencies in time series data.
DLinear operates by decomposing the time series into a trend component and a remainder component. It then applies two separate one-layer linear networks to model these components independently. This direct multi-step (DMS) forecasting approach contrasts with the autoregressive methods commonly used in traditional models, which can suffer from error accumulation over long prediction horizons.
Despite its simplicity, DLinear has demonstrated superior performance across various benchmarks, often surpassing Transformer-based models. Its strengths lie in its ability to capture both short-term and long-term dependencies efficiently, with a low computational cost and high interpretability. DLinear’s success raises important questions about the necessity of using complex Transformer architectures for LTSF and suggests a reevaluation of their application in time series forecasting tasks.
A.1.4 NSTformer [6]
@@ A.2 Datasets
Weather dataset The Weather dataset, provided by the Max Planck Institute for Biogeochemistry, comprises detailed meteorological measurements from 2020 to 2021. Recorded every ten minutes, this dataset includes variables such as temperature, humidity, and wind speed. It is essential for high-resolution weather forecasting and climate research, offering insights into microclimatic fluctuations and seasonal patterns. This rich dataset serves as a valuable resource for understanding and predicting weather-related changes over short intervals.
Table 4: Summary of Datasets Used in Time Series Forecasting
@@
Weather 52695 10Min 21
A.3 Definition of Metric
@@
$\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}\left|y_{i}-\hat{y}_{i}\right|$ (14)
@@
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}-\hat{y}_{i}\right)^{2}$ (15)
@@ A.4 Experimental Setup
Table 5: Experimental Setup: Summary of Hardware and Software Configurations
Appendix B Comparison of of different $\mathcal{L}_{IB}$ loss weights
Table 6: Comparison of different $\mathcal{L}_{IB}$ loss weights on Exchange rate dataset. The lookback length is set to $T=2H$.
@@ Appendix B Comparison of of different $\mathcal{L}_{IB}$ loss weights
48 0.1394 0.0505 0.1391 0.0451 0.1390 0.0448 0.1389 0.0441 0.1390 0.0440
96 0.1947 0.1102 0.1950 0.0833 0.1946 0.0834 0.1958 0.0857 0.1950 0.0854
144 0.2491 0.1392 0.2414 0.1333 0.2413 0.1133 0.2411 0.1535 0.2403 0.1527
192 0.3008 0.1886 0.2840 0.1802 0.2836 0.1803 0.2834 0.1778 0.2893 0.1804
Table 6 presents the impact of varying information bottleneck ($\mathcal{L}_{IB}$) loss weights on model performance across different forecast horizons on the Exchange rate dataset. Notably, at a 48-step forecast, the lowest MAE and MSE were observed at IB weights of 0.01 and 0.1, respectively, suggesting that moderate IB weights help minimize predictive errors while maintaining accuracy.
At a 96-step forecast, the optimal MAE appeared at an IB weight of 0.001, while the lowest MSE was at 0.00001, indicating that smaller IB weights can sustain lower errors over longer forecasts.
For longer forecasts (144 and 192 steps), similar trends were observed with the lowest MAE and MSE appearing at adjusted IB weights, indicating that the selection of appropriate IB weights significantly influences model performance. However, when the IB weight is reduced beyond a certain threshold, the performance improvements plateau, suggesting that overly small IB weights might lose their effectiveness.
Appendix C Comparison of wavelet basis functions
Table 7: Comparison of different wavelet basis functions on Exchange rate dataset. The lookback length is set to $T=2H$.
@@ Appendix C Comparison of wavelet basis functions
48 0.1392 0.0478 0.1420 0.0450 0.1390 0.0448
96 0.1945 0.1024 0.1945 0.0895 0.1946 0.0834
144 0.2446 0.1230 0.2430 0.1326 0.2413 0.1133
192 0.3007 0.2840 0.2914 0.1804 0.2840 0.1803
Table 7showcases the comparative performance of different wavelet basis functions—HAAR, SYM2, and DB1—on the Exchange rate dataset with a lookback length set to $T=2H$. The performance metrics used are MAE and MSE, which provide insights into the accuracy and reliability of the predictions at various forecast horizons.
For a 48-step forecast, DB1 demonstrates the lowest MAE and MSE, indicating its superior ability to capture the underlying patterns in the data at shorter horizons. HAAR, while slightly less effective than DB1 in terms of MSE, offers comparable MAE values, suggesting that it remains a robust choice for certain types of financial data.
At a 96-step forecast, HAAR shows the lowest MAE, matching closely with SYM2, but DB1 outperforms both in terms of MSE, highlighting its efficiency in reducing prediction errors over longer horizons.
For 144 and 192-step forecasts, DB1 consistently shows the best performance in both MAE and MSE, underlining its effectiveness in handling longer-term dependencies in the dataset. SYM2, while competitive, tends to exhibit slightly higher errors, particularly in MSE, suggesting that its utility may be more limited under conditions requiring high precision.
Overall, the selection of the appropriate wavelet basis function significantly impacts model performance, with DB1 generally offering the best balance of accuracy and error reduction across the board. This analysis underscores the importance of choosing the right wavelet transform to enhance predictive performance in time series forecasting, especially in the context of financial datasets.